This is exactly how it is typed:
A number is 6 greater than half another number. If the sum of the numbers is 21, find the numbers.

Found 2 solutions by richwmiller, josgarithmetic:
Answer by richwmiller(17219)

a=b/2+6
a+b=21
21-b=b/2+6
15=3/2b
30=3b
10=b
a=11


Answer by josgarithmetic(39620)

The description is ambiguous. The first part seems to be saying, 6+(1/2)n, if one of the numbers is m and the other is n.

Letting m and n be the numbers,
m=6+(1/2)n and m+n=21.
Substitute for m in the second equation.
(6+(1/2)n)+n=21
6+n/2+n=21
(3/2)n=21-6
(3/2)n=15
n=(2/3)15
n=10 and this means m=11.